What mass of AgCl(s) is formed when a solution containing 2.235 g of KCl is mixed with a solution contains 7.65 g of AgNO3 ?
(Ar(ag)=108;Ar(CL)=35.5)
1
Expert's answer
2020-02-14T07:10:53-0500

KCl+AgNO3=AgCl+KNO3

Find the amount of substance in moles

v=m/M

v(KCl)=2.235/(35.453+39.098)=2.235/74.551=0.029 mol

v(AgNO3)=7.65/(107.868+15.999*3+14.0067)=7.65/169.8717=0.045 mol


The reaction equation shows that n (AgNO3): n (KCl) = 1: 1

then KCl is in shortage, and AgNO3 is in excess. We further calculate KCl:

v(AgCl)=0.029

m=v*M=0.029*(108+35.5)=0.029*143.5=4.162 g